.- Let [2 2 11 A = 12 5 2 , |1 2 2 | fіnd det(A – XI) = 0
Answer #1

Given that 2 2 17 A = 2 5 2 To show, det ( -)I) -O consider, A-AT (2217 sioon 52- LI 221 Lo - [ 2 21 (A 0 | 25 2-lono Li ? 202b-keltzk8-EL : 0=&X+218-5-K21 Ö 0 = 5 Et qt-ottis x3_8x²+ 137-6=0 which is required solution,